A box contains 5 red, 4 blue and 3 white balls. In how many ways can we select 3 balls such that
a. They are of different colors?
b. They are all red?
c. Two are blue and one is white?
d. None is blue?













Answer by sudhanshu_kmr(1152)


a) here one ball of each color

no. of ways = 5C1 * 4C1 * 3C1 = 60

b) no. of ways = 5C3 = 10

c) no. of ways = 4C2 * 3C1 = 6 * 3 = 18

d) number of non-blue ball = 8
no. of ways = 8C3 = 56